# Mr. DETR: Instructive Multi-Route Training for Detection Transformers

### v1

@@ Abstract
Existing methods enhance the training of detection transformers by incorporating an auxiliary one-to-many assignment. In this work, we treat the model as a multi-task framework, simultaneously performing one-to-one and one-to-many predictions. We investigate the roles of each component in the transformer decoder across these two training targets, including self-attention, cross-attention, and feed-forward network. Our empirical results demonstrate that any independent component in the decoder can effectively learn both targets simultaneously, even when other components are shared. This finding leads us to propose a multi-route training mechanism, featuring a primary route for one-to-one prediction and two auxiliary training routes for one-to-many prediction. We enhance the training mechanism with a novel instructive self-attention that dynamically and flexibly guides object queries for one-to-many prediction. The auxiliary routes are removed during inference, ensuring no impact on model architecture or inference cost. We conduct extensive experiments on various baselines, achieving consistent improvements as shown in Fig. 1.
@@ 1 Introduction
The end-to-end detection transformer (DETR) [3], along with subsequent studies [33, 40, 50], shine in object detection due to its simplicity and effectiveness. Unlike traditional methods, DETR-based detectors eliminate the need for post-processing steps like non-maximum-suppression (NMS). They employ the one-to-one assignment strategy (see Fig. 2(a)) for supervision, where each ground-truth box is matched to a single prediction. Compared to the one-to-many assignment methods (see Fig. 2(b)) in conventional object detectors [43, 32, 49] that allow each ground-truth box to be assigned with multiple predictions, the one-to-one assignment can lead to slow convergence [20, 58, 47] of DETRs due to its sparse supervision.
@@ 3.1 Preliminary
Traditional object detectors [43, 32, 49] typically use a one-to-many assignment strategy, assigning each ground truth box to multiple predictions based on specific criteria, followed by NMS to eliminate duplicated predictions. In our work, we apply a straightforward one-to-many assignment strategy [41] as in [62, 19], which considers localization quality and classification confidence, making it suitable for DETR-like detectors. Specifically, the matching score $M_{ij}$ between the predictions $(s_{i},b_{i})$ and the ground truth $(\bar{s_{t}},\bar{b_{t}})$ is defined as:
@@ 3.3 Instructive Self-Attention
Specifically, we build $m$ learnable tokens $\mathbf{Q^{ins}}=\{q_{0}^{ins},q_{1}^{ins},...,q_{m-1}^{ins}\}$, called instruction tokens. Initially, these instruction tokens are attached to the input sequence of self-attention by concatenation, forming a composite set of input queries $\mathbf{\hat{Q}^{ins}}=\{q_{0}^{ins},q_{1}^{ins},...,q_{m-1}^{ins},q_{0},q_{1},...,q_{n-1}\}$, resulting in a length of $m+n$. Subsequently, self-attention is performed on these combined queries. As the attached instruction tokens are not utilized for object localization, their outputs are discarded post self-attention, serving solely to convey information to object queries. The output $\hat{\mathbf{Q}}_{3}$ of Route-3 is written as:
@@
Discussion.  Large language models have demonstrated impressive capabilities [39] when equipped with meticulously crafted prompts [46, 22]. Subsequent researches treat these prompts as learnable tokens, achieving success in both natural language processing [28, 25, 35] and computer vision applications [66, 23, 51]. While prompt tuning [21, 45, 55] and our proposed instructive self-attention both involve attaching learnable tokens to the input sequence, our instruction token uniquely facilitates distinguishing between one-to-one and one-to-many assignments, rather than merely fine-tuning the model.
@@
Conditional-DETR [40] 50 300 40.9 61.8 43.3 20.8 44.6 60.2
Anchor-DETR [50] 50 300 42.1 63.1 44.9 22.3 46.2 60.0
SAM-DETR [56] 50 300 39.8 61.8 41.6 20.5 43.4 59.6
DAB-DETR [33] 12 300 44.2 62.5 47.3 27.5 47.1 58.6
DN-DETR [26] 12 300 46.0 63.8 49.9 27.7 49.1 62.3
SQR-DETR [4] 50 300 45.9 64.7 50.2 27.7 49.2 60.5
@@ 4.1 Setup
Dataset and Evaluation. The COCO 2017 dataset [31], widely utilized for object detection and instance segmentation, serves as the benchmark for our model evaluation. In line with existing research [67, 58, 20], we conduct evaluations on the COCO 2017 validation set, providing results in terms of standardized metrics, namely average precision (mAP, AP50, AP75) at various IoU thresholds. APs, APm, and APl refer to the average precision for small, medium, and large objects, respectively. For all experiments with 300 and 900 queries, we evaluate the performance according to top-100 and top-300 predictions, respectively.
@@
Rank-DETR [42] 12 900 57.6 76.0 63.4 41.6 61.4 73.8
DAC-DETR [19] 12 900 57.3 75.7 62.7 40.1 61.5 74.4
EASE-DETR [13] 12 900 57.8 76.7 63.3 40.7 61.9 73.7
@@
Table 3: The performance on the COCO 2017 [31] validation set. All models are based on the Swin-L [37] backbone.
@@ 4.2 Main Results
We conduct extensive experiments on the COCO 2017 dataset. In Tab. 2, we present the performance of our method across various baselines, including Deformable-DETR++ [67] with either 300 or 900 queries, DINO [58] and Align-DETR [1]. Specifically, using Deformable-DETR++ [67] with 300 queries, our method achieves a 49.5% mAP, providing a 2.5% improvement without additional inference cost. It surpasses other variants, such as H-DETR [20], MS-DETR [62] and DAC-DETR [19], by approximately 0.7% in mAP. Based on a strong baseline model like Deformable-DETR++ [67] with 900 queries, our multi-route training method reaches a 50.7% mAP under a 12-epoch training schedule. Extending to 24 epochs, our approach enables Deformable-DETR++ [67] to achieve a 51.4% mAP, outperforming the baseline by 1.6%. Compared to other variants with one-to-many auxiliary training, our method exceeds DAC-DETR [19] and MS-DETR [62] by 0.9% and 0.5%, respectively.
Using our approach, DINO baseline [58] can reach 50.9% mAP over 12 epochs, outperforming variants like Group-DETR [5], DAC-DETR [19] and MS-DETR [62]. Previous studies [54, 13, 1, 42, 34, 18] enhance model architectures or implement IoU-aware losses, such as VFL loss [57] and IA-BCE loss [1]. To demonstrate our method’s compatibility with these advancements, we apply it to Align-DETR [1], achieving gains of 1.2% and 1.0% mAP in 12 and 24 epoch schedules, respectively. Our experiments consistently demonstrate improvements across different baseline models. Further experiments with the Swin-L [37] backbone, as shown in Tab. 3, reveal that our method achieves a 58.4% mAP in a 12-epoch training schedule, surpassing other variants. These experimental results underscore the effectiveness of our proposed method.
@@ 4.5 Ablation Study
Our method features a primary route, referred to as Route-2 in Fig. 4, dedicated to one-to-one prediction, and two auxiliary routes, Route-1 and Route-3 for one-to-many prediction. Route-1 utilizes shared self-attention and cross-attention with Route-2, but employs an independent FFN to achieve one-to-many prediction. We introduce an instructive self-attention mechanism in Route-3, which shares all parameters with the primary route. To evaluate the contribution of each auxiliary route, we perform an ablation study, as shown in Tab. 5. Route-1, with its independent FFN, enhances the baseline model by 2.0% mAP. Experimental results indicate that Route-3 provides a 2.8% improvement on Deformable-DETR++ [67], highlighting the effectiveness of our proposed instructive self-attention. By incorporating multi-route training, Deformable-DETR++ [67] achieves 50.7% mAP, surpassing the baseline by 3.1%.
@@ Acknowledgement
This work is supported by Hong Kong Research Grant Council - Early Career Scheme (Grant No. 27208022), National Natural Science Foundation of China (Grant No. 62306251), and HKU Seed Fund for Basic Research. The computations were performed partly using research computing facilities offered by Information Technology Services, the University of Hong Kong.
Appendix A Pseudo-Code for Instructive Self-Attention
@@
Algorithm 1 Pseudo-code of Pooled Cube Distillation in a PyTorch-like style.
@@
query = torch.cat([ins_token + ins_pos, query], dim=1)
key = torch.cat([ins_token + ins_pos, key], dim=1)
value = torch.cat([ins_token, value], dim=1)
@@
Appendix B Convergence Curves
(a) Evaluation results of box
(b) Evaluation results of mask
Appendix C Detailed Performance of Each Route
As detailed in Sec. B, we train Mr. DETR for the instance segmentation task based on Deformable-DETR++ [67]. We report the evaluation results of each route of our method in Tab. 8. Experimental results indicate that the primary route is adept at accomplishing one-to-one prediction, as evidenced by the fact that Non-Maximum-Suppression (NMS) yields only a minor improvement of approximately 0.1 - 0.2% in both box mAP and mask mAP. Conversely, for the auxiliary routes, the application of NMS significantly improves the performance (about 35% and 25% in terms of box mAP and mask mAP) of Route-1 and Route-3, highlighting their capability for effective one-to-many prediction. This further substantiates that our introduced instructive self-attention is proficient in efficiently guiding object queries for one-to-many prediction.
@@ Appendix C Detailed Performance of Each Route
Appendix D Detailed Performance of Intermediate Layers
Typically, DETR-like object detectors consist of six layers each in their transformer encoders and decoders. As mentioned in Sec. B and Sec. C, our approach for mask prediction is based on Deformable-DETR++ [67], utilizing solely the last decoder layer. All six decoder layers are employed for object detection tasks. Therefore, we only evaluate the box prediction for all layers as shown in Tab. 9. Evaluation results suggest that our method can effectively improve the performance of the primary route across all six decoder layers, demonstrating the efficacy of our approach. Moreover, the one-to-many prediction training routes, namely, Route-1 and Route-3, significantly surpass the primary route in the shallower layers. For example, with a model trained on a 12-epoch schedule, Route-3 achieves a 6.4% improvement over the primary route in layer 0, and a 0.4% improvement in layer 5. These experiments indicate that the primary route needs more decoder layers to reach comparable performance as the auxiliary routes equipped with NMS. For instance, Route-1 and Route-3 can reach 49.4% mAP in layer 2, while the primary route achieves 49.4% mAP in layer 4.
@@ Appendix D Detailed Performance of Intermediate Layers
Appendix E Training Cost
We measure the training time of different methods in Tab. 10. The training time denotes the average duration of each epoch. We evaluate training time on 8 NVIDIA 3090 GPUs with a batch size of 16. The experimental results indicate that our proposed method achieves an effective trade-off between performance and training time.
Model Training Time (minutes) mAP
Deformable-DETR++ [67] 84 47.0
H-DETR [20] 104 (+20) 48.7 (+1.7)
DAC-DETR [19] 94 (+10) -
MS-DETR [62] 96 (+12) 48.8 (+1.8)
Group-DETR [5] 123 (+39) -
Mr. DETR (Ours) 101 (+17) 49.5 (+2.5)
Table 10: Comparison of training time of various methods. All methods utilize Deformable-DETR++ with 300 queries as the baseline. The training time represents the average duration per training epoch.
Appendix F Qualitative Results
We present the prediction results of our method in Fig. 8. The model is based on the DINO [58] baseline with Swin-L [37] backbone.
Figure 8: Qualitative results of our method. Left: prediction results. Right: ground-truth.
Appendix G Impact of Hyper-Parameters
(a) Ablation of $K$
(b) Ablation of $\alpha$
(c) Ablation of $\tau$

### v3

@@ Abstract
Existing methods enhance the training of detection transformers by incorporating an auxiliary one-to-many assignment. In this work, we treat the model as a multi-task framework, simultaneously performing one-to-one and one-to-many predictions. We investigate the roles of each component in the transformer decoder across these two training targets, including self-attention, cross-attention, and feed-forward network. Our empirical results demonstrate that any independent component in the decoder can effectively learn both targets simultaneously, even when other components are shared. This finding leads us to propose a multi-route training mechanism, featuring a primary route for one-to-one prediction and two auxiliary training routes for one-to-many prediction. We enhance the training mechanism with a novel instructive self-attention that dynamically and flexibly guides object queries for one-to-many prediction. The auxiliary routes are removed during inference, ensuring no impact on model architecture or inference cost. We conduct extensive experiments on various baselines, achieving consistent improvements as shown in Fig. 1. Project page: https://visual-ai.github.io/mrdetr
@@ 1 Introduction
The end-to-end detection transformer (DETR) [3], along with subsequent studies [34, 41, 53], shine in object detection. Unlike traditional methods, DETR-based detectors eliminate the need for non-maximum-suppression (NMS). They employ the one-to-one assignment strategy (see Fig. 2(a)) for supervision, where each ground-truth box is matched to a single prediction. Compared to the one-to-many assignment methods (see Fig. 2(b)) in conventional object detectors [44, 33, 52] that allow each ground-truth box to be assigned with multiple predictions, the one-to-one assignment can lead to slow convergence [21, 61, 50] of DETRs due to its sparse supervision.
@@ 3.1 Preliminary
Traditional object detectors [44, 33, 52] typically use a one-to-many assignment strategy, assigning each ground-truth box to multiple predictions based on specific criteria, followed by NMS to eliminate duplicated predictions. In our work, we apply a straightforward one-to-many assignment strategy [42] as in [65, 20], which considers localization quality and classification confidence, making it suitable for DETR-like detectors. Specifically, the matching score $M_{ij}$ between the predictions $(s_{i},b_{i})$ and the ground truth $(\bar{s_{t}},\bar{b_{t}})$ is defined as:
@@ 3.3 Instructive Self-Attention
Specifically, we build $m$ learnable tokens $\mathbf{Q^{ins}}=\{q_{0}^{ins},q_{1}^{ins},...,q_{m-1}^{ins}\}$, called instruction tokens. Initially, these instruction tokens are attached to the input sequence of self-attention by concatenation, forming a composite set of input queries $\mathbf{\hat{Q}^{ins}}=\{q_{0},q_{1},...,q_{n-1},q_{0}^{ins},q_{1}^{ins},...,q_{m-1}^{ins}\}$, resulting in a length of $n+m$. Subsequently, self-attention is performed on these combined queries. As the attached instruction tokens are not utilized for object localization, their outputs are discarded post self-attention, serving solely to convey information to object queries. The output $\hat{\mathbf{Q}}_{3}$ of Route-3 is written as:
@@
Discussion.  Large language models have demonstrated impressive capabilities [40] when equipped with meticulously crafted prompts [49, 23]. Subsequent researches treat these prompts as learnable tokens, achieving success in both natural language processing [29, 26, 36] and computer vision applications [69, 24, 54]. While prompt tuning [22, 48, 58] and our proposed instructive self-attention both involve attaching learnable tokens to the input sequence, our instruction token uniquely facilitates distinguishing between one-to-one and one-to-many assignments, rather than merely fine-tuning the model. Additionally, different from the typical Mixture-of-Experts methods [47, 11] that dynamically select experts based on the input feature through a gating mechanism, our separated FFN pre-assigns an expert to each specific task.
@@ 4.1 Setup
Dataset and Evaluation. In line with existing research [70, 61, 21], we conduct experiments on the COCO 2017 [32], providing results in terms of standardized metrics, namely average precision (mAP, AP50, AP75) at various IoU thresholds. APs, APm, and APl refer to the average precision for small, medium, and large objects, respectively. For all experiments with 300 and 900 queries, we report performance according to top-100 and top-300 predictions, respectively.
@@
Rank-DINO [43] 12 900 57.6 76.0 63.4 41.6 61.4 73.8
DAC-DINO [20] 12 900 57.3 75.7 62.7 40.1 61.5 74.4
EASE-DINO [14] 12 900 57.8 76.7 63.3 40.7 61.9 73.7
@@
Table 3: The performance on the COCO 2017 [32] validation set. All models use the Swin-L [38] as the architecture backbone. In this experiment, Mr. DETR is built upon the DINO detection framework [61] with IA-BCE loss [1]. The denotes the results derived from [35].
@@ 4.2 Main Results
We conduct extensive experiments on the COCO 2017 dataset [32]. As shown in Tab. 2, in the training schedule of 12 epochs, our method achieves consistent improvement across various baseline models. Extending to 24 epochs, our approach enables Deformable-DETR++ [70] to achieve a 51.4% mAP, outperforming the baseline by 1.6%. Compared to other variants with one-to-many auxiliary training, our method exceeds DAC-DETR [20] and MS-DETR [65] by 0.9% and 0.5%, respectively. Using our approach, DINO baseline [61] can reach 50.9% mAP over 12 epochs, outperforming variants like Group-DETR [5], DAC-DETR [20] and MS-DETR [65]. To demonstrate our method’s compatibility with localization-aware loss [60, 1], we apply it to Align-DETR [1], achieving gains of 1.2% and 1.0% mAP in 12 and 24 epoch schedules, respectively. Our experiments consistently demonstrate improvements across different baseline models. Further experiments with the Swin-L [38] backbone, as shown in Tab. 3, reveal that our method achieves a 58.4% mAP in a 12-epoch training schedule, surpassing other variants. These experimental results underscore the effectiveness of our proposed method.
@@ 4.5 Ablation Study
To evaluate the contribution of each auxiliary route, we perform an ablation study, as shown in Tab. 5. Route-1, with its independent FFN, enhances the baseline model by 2.0% mAP. Experimental results indicate that Route-3 provides a 2.8% improvement on Deformable-DETR++ [70], highlighting the effectiveness of our proposed instructive self-attention. By incorporating multi-route training, Deformable-DETR++ [70] achieves 50.7% mAP, surpassing the baseline by 3.1%.
@@ Acknowledgement
This work is supported by National Natural Science Foundation of China (Grant No. 62306251), Hong Kong Research Grant Council - Early Career Scheme (Grant No. 27208022), and HKU Seed Fund for Basic Research. The computations were performed partly using research computing facilities offered by Information Technology Services, The University of Hong Kong.
Appendix
Appendix A Analysis of Instructive Self-Attention
Figure 7: The cosine similarity between 10 instruction tokens used in Deformable-DETR++ [70].
@@ Appendix A Analysis of Instructive Self-Attention
(a) Evaluation results of box
(b) Evaluation results of mask
Algorithm 1 Pseudo-code of Instructive Self-Attention in a PyTorch-like style.
@@
query = torch.cat([query, ins_token + ins_pos], dim=1)
key = torch.cat([key, ins_token + ins_pos], dim=1)
value = torch.cat([value, ins_token], dim=1)
@@
(a) Ablation of $K$
(b) Ablation of $\alpha$
(c) Ablation of $\tau$
In the main paper, we ablate the configurations of the proposed instructive self-attention, including the number of instruction tokens and layers using instruction tokens. Ablation studies demonstrate that the model is not sensitive to the layers and numbers of instruction tokens, due to the following reasons: (i) Regarding the impact of layers of instruction tokens, we hypothesize that the information from instruction tokens in the first layer can be retained and utilized by subsequent layers. Meanwhile, residual connections across transformer decoder layers may help preserve the instruction information for later layers. (ii) To study the impact of the number of instruction tokens, as shown in Fig. 7, we calculate the cosine similarity between the 10 instruction tokens and find that most of them are very similar. This indicates that most instruction tokens may play a similar role, making the model insensitive to their number.
@@
Appendix B Convergence Curves
Appendix C Impact of Hyper-Parameters
Appendix D Detailed Performance of Each Route
As detailed in Sec. B, we train the Mr. DETR for the instance segmentation task based on Deformable-DETR++ [70]. We report the evaluation results of each route of our method in Tab. 8. Experimental results indicate that the primary route is adept at accomplishing one-to-one prediction, as evidenced by the fact that Non-Maximum-Suppression (NMS) yields only a minor improvement of approximately 0.1 - 0.2% in both box mAP and mask mAP. Conversely, for the auxiliary routes, the application of NMS significantly improves the performance (about 35% and 25% in terms of box mAP and mask mAP) of Route-1 and Route-3, highlighting their capability for effective one-to-many prediction. This further substantiates that our introduced instructive self-attention is proficient in efficiently guiding object queries for one-to-many prediction.
Models Epochs Queries mAP AP50 AP75 APs APm APl
Deformable-DETR++ [70] 4 900 30.4 40.8 33.1 16.1 30.1 39.1
w/ Mr. DETR 4 900 32.7 (+2.3) 42.7 (+1.9) 35.8 (+2.7) 17.1 (+1.0) 32.3 (+2.2) 42.6 (+3.5)
Table 9: Experiments on the large-scale Obejcts365 dataset [46]. Deformable-DETR++ and our model use the ResNet-50 [17] backbone.
Appendix E Experiments on the Objects365
Objects365 [46] is a large-scale dataset with 365 classes, which contains about 2,000,000 images. To further verify the scalability of our method on the large-scale dataset, we conduct experiments on Objects365 using Deformable-DETR++ [70] model with 900 queries. To save training time, we train the baseline model and our method for 4 epochs only. The initial learning rate is set to 2e-4 and decays at the third epoch. Other training settings are the same as the model trained on the COCO dataset [32]. We report the performance comparison in Tab. 9.
Appendix F Training Cost
We measure the training time of different methods in Tab. 10. The training time denotes the average duration of each epoch. We evaluate training time on 8 NVIDIA 3090 GPUs with a batch size of 16. The experimental results indicate that our proposed method achieves an effective trade-off between performance and training costs.
Model Time GFLOPs mAP
Deformable-DETR++ [70] 84 234.9 47.0
H-DETR [21] 104 (+20) 265.0 48.7 (+1.7)
DAC-DETR [20] 94 (+10) - -
MS-DETR [65] 96 (+12) - 48.8 (+1.8)
Group-DETR [5] 123 (+39) - -
Mr. DETR (Ours) 101 (+17) 258.0 49.5 (+2.5)
Table 10: Comparison of training time (minutes) and training GFLOPs of various methods. All methods utilize Deformable-DETR++ with 300 queries as the baseline. The training time represents the average duration per training epoch.
Appendix G Performance of Intermediate Layers
Typically, DETR-like object detectors consist of six layers each in their transformer encoders and decoders. As mentioned in Sec. B and Sec. D, our approach for mask prediction is based on Deformable-DETR++ [70], utilizing solely the last decoder layer. All six decoder layers are employed for object detection tasks. Therefore, we only evaluate the box prediction for all layers as shown in Tab. 11. Evaluation results suggest that our method can effectively improve the performance of the primary route across all six decoder layers, demonstrating the efficacy of our approach. Moreover, the one-to-many prediction training routes, namely, Route-1 and Route-3, significantly surpass the primary route in the shallower layers. For example, with a model trained on a 12-epoch schedule, Route-3 achieves a 6.4% improvement over the primary route in layer 0, and a 0.4% improvement in layer 5. These experiments indicate that the primary route needs more decoder layers to reach comparable performance as the auxiliary routes equipped with NMS. For instance, Route-1 and Route-3 can reach 49.4% mAP in layer 2, while the primary route achieves 49.4% mAP in layer 4.
@@ Appendix G Performance of Intermediate Layers
Appendix H Qualitative Results
We present the prediction results of our method in Fig. 10. The model is based on the DINO [61] with IA-BCE loss [1] as the baseline using Swin-L [38] backbone.
Figure 10: Qualitative results of our method. Left: prediction results. Right: ground truth.